on a right triangle if a=4 and c=2radical13 what is b?
Answer by rfadrogane(214)

on a right triangle if a=4 and c=2√13, what is b?
Sol,n:
c^2 = a^2+b^2
b^2 = c^2 - a^2
b^2 = (2√13)^2 - (4)^2
b = √[52-16]
b = 6